OpenStudy (anonymous):

if the half-life of a unstable isotope is 10,00 years and only 1/8 of the radioactive parent remains how old is the sample? 10,000 20,000 30,000 or 40,000

thomaster (thomaster):

\(\Huge\color{gold}{\star\star\star\star\star\star\hspace{7pt}}\LARGE\tt\color{orangered}{Hey\;there,}\hspace{2pt}\Huge\color{gold}{\star\star\star\star\star\star\star\star}\\\Huge\hspace{15pt}\tt\Huge\;\color{Blue}{welcome\;to\;openstudy!}\\\Huge\color{gold}{\star\hspace{1pt}\star\star\star\star\star\star\star\star\star\star\star\star\star\star\star\star\star\star\star\star}\) The half life is 10,000 years i assume? This is the time when \(\frac{1}{2}\) of the isotopes is left. After 2 half lifes there is \(\frac{1}{2}*\frac{1}{2}=\frac{1}{4}\) left After 3 half lifes there is \(\frac{1}{4}*\frac{1}{2}=\frac{1}{8}\) left There is \(\frac{1}{8}\) left after 3 half lifes, which is 3 times 10,000 years. So the answer is 30,000 years :D \(\overline{\underline{\LARGE{\color{gold}{\star~}}\Large\tt\color{green}{I\;Hope\;this\;Helps!}\LARGE{\color{gold}{~\star}}}}\)